# Theories of synaptic memory consolidation and intelligent plasticity for continual learning

### v1

@@ 5 Conclusion
Finally, we must progress in connecting our top-down motivated models that have primarily been derived using a deep learning framework to phenomenological plasticity models and realistic spiking neural network models. This step will allow interfacing theory and experiment and is thus essential to test whether and how neurobiology leverages similar continual learning principles. In particular, we have to work towards integrative theories that combine insights from neurobiology, such as cross-tagging, consolidation, and reconsolidation (Alberini, 2005), which presently do not have an apparent correspondence in the functional continual learning models developed for deep neural networks. Studying these phenomena may also shed light on why humans and animals are typically better at blocked training (Flesch et al., 2018), whereas ANNs still excel at joint training. To address these questions, we must derive testable predictions from the models and systematically verify them through experiments. Doing so will likely require a multi-scale approach in which we derive simultaneous predictions at the network, microcircuit, neuronal, and synaptic levels. Models that accurately capture continual learning in the brain must also accurately account for effects across scales.

### v2

@@ 5 Conclusion
Finally, we must progress in connecting our top-down motivated models that have primarily been derived using a deep learning framework to phenomenological plasticity models and realistic spiking neural network models. This step will allow interfacing theory and experiment and is thus essential to test whether and how neurobiology leverages similar continual learning principles. In particular, we have to work towards integrative theories that combine insights from neurobiology, such as cross-tagging, consolidation, and reconsolidation (Alberini, 2005; Gershman et al., 2017), which presently do not have an apparent correspondence in the functional continual learning models developed for deep neural networks. Studying these phenomena may also shed light on why humans and animals are typically better at blocked training (Flesch et al., 2018), whereas ANNs still excel at joint training. To address these questions, we must derive testable predictions from the models and systematically verify them through experiments. Doing so will likely require a multi-scale approach in which we derive simultaneous predictions at the network, microcircuit, neuronal, and synaptic levels. Models that accurately capture continual learning in the brain must also accurately account for effects across scales.